# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ Abstract
Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, yet their robustness in code comprehension and reasoning remains insufficiently explored. We present CodeCrash, a comprehensive stress-testing benchmark comprising 1,279 questions from two established datasets, CruxEval and LiveCodeBench, designed to evaluate model reasoning reliability under non-standard coding environments. We systematically evaluate 17 LLMs across input and output prediction tasks using direct and Chain-of-Thought prompting approaches, revealing that LLMs are particularly vulnerable to disorganized code and overly reliant on natural language cues: aggregated structural perturbations result in over 14 percentage points (pp) of degradation, while textual perturbations cause a performance drop of over 11 pp. Moreover, self-reflective mechanisms in state-of-the-art reasoning models significantly increase token usage by 23×2-3$\times$2 - 3 ×, reduce output confidence, and even lead to catastrophic reasoning failures when faced with targeted perturbations—for instance, QwQ-32B generates over 12,000 redundant tokens under reasoning-level perturbations. CodeCrash provides a rigorous benchmark for evaluating robustness in code understanding, guiding future research toward more reliable and resilient LLMs in code reasoning. The benchmark code, perturbed datasets, and full leaderboard are publicly available at https://cuhk-arise.github.io/CodeCrash/.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by GPT families OpenAI et al. (2024); Hurst et al. (2024) and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Liu et al. (2023b); Xiao et al. (2024), code completion Ding et al. (2023), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). These advancements have facilitated the incorporation of LLMs into Integrated Development Environments, such as GitHub Copilot GitHub (2022), OpenAI Codex Chen et al. (2021), and Tabnine Tabnine (2023), highlighting their potential to automate various aspects of software development Zhang et al. (2023); Pandey et al. (2024).
While the effectiveness of these applications heavily depends on the model’s ability to comprehend the intended program logic, underlying functionality, and developer intent Pan et al. (2024), real-world production codebases are often disorganized, with ambiguous identifiers Gao et al. (2023) and inconsistent comments Rong et al. (2025). Such non-standard artifacts can mislead models, leading to incorrect inferences and misinterpretations, compromising the model’s reliability in downstream tasks. Recent studies Wang et al. (2022); Mastropaolo et al. (2023); Chen et al. (2024a) have primarily evaluated LLM robustness under prompt-level perturbation, focusing on variations in task descriptions. While insightful for prompt engineering, these studies overlook robustness in deeper code comprehension tasks, such as reasoning about execution logic and program semantics.
To determine whether LLMs truly understand program logic, we adopt input and output prediction tasks introduced in CruxEval (CruxGu et al. (2024) as diagnostic methods for assessing model in code understanding and execution reasoning. Inspired by prior work on stress-testing Natural Language (NL) inference systems Naik et al. (2018), we propose a novel stress-testing strategy to evaluate the robustness of LLMs’ code comprehension capabilities under systematic code perturbations. First, building upon program structure-consistent mutations Li et al. (2023) and dead-loop poisoning Wan et al. (2022); Sun et al. (2023), we design Code Structural Perturbations, which involve logic-preserving transformations, such as renaming entities, reformatting conditional expressions, and inserting unreachable dead code. Second, motivated by inconsistent comments in real-world codebases Rong et al. (2025), we introduce Textual Distraction Perturbations, which inject misleading NL information into code to challenge the model’s ability to interpret program logic under conflicting cues.
To extend our evaluation from synthetic programs to real-world codebases, we incorporate LiveCodeBench (LCBJain et al. (2025) alongside Crux as part of our benchmark suite. We evaluate seventeen LLMs across distinct perturbations in input and output prediction tasks. Additionally, we apply Chain-of-Thought (CoT) prompting Kojima et al. (2022) to explore how explicit step-by-step reasoning enhances model performance and robustness in code comprehension. Furthermore, we assess three state-of-the-art Large Reasoning Models (LRMs) with powerful internal reasoning processes, uncovering significant vulnerabilities arising from their self-reflection mechanisms. Our key findings reveal that:
  • LLMs perform unreliably under perturbations and heavily rely on embedded NL cues for reasoning and comprehension. Each textual distraction perturbation causes an average performance drop of over $11$pp 111pp stands for percentage point., while the aggregated structural perturbation also leads to up to $14$pp degradation.
  • CoT prompting improves model robustness, but cannot fully eliminate sensitivity to NL cues, resulting in a relative degradation of up to $14\%$ under contextual-level textual perturbation.
  • LRMs show superior robustness under most perturbations, but the self-reflection mechanism consumes 23×2-3$\times$2 - 3 × more reasoning tokens to resolve contradictions, even leading to reflection loops.
@@
  • We introduce a comprehensive perturbation framework, CodeCrash, to evaluate the code-level robustness of LLMs, which is easily applicable to other benchmarks and programming languages.
  • We conduct extensive experiments on 17 LLMs across two established datasets, Crux and LCB.
  • We provide novel insights into evaluating and interpreting code reasoning in LLMs, revealing the reliance on NL cues and demonstrating how reasoning-level hints can lead to model collapse.
2 Related Work
Researchers have significantly contributed to developing benchmarks for evaluating LLM capabilities on code-related tasks. Early benchmarks such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021), APPS Hendrycks et al. (2021), and HumanEval+ Liu et al. (2023b) primarily focused on assessing code generation performance. More recent efforts have shifted toward practical software engineering and realistic coding scenarios. For instance, SWE-bench Jimenez et al. (2024), InfiBench Li et al. (2024), and LCB Jain et al. (2025) evaluate LLMs on complex tasks like multi-file changes, Stack Overflow-style questions, and real-world programming problems.
Meanwhile, a growing number of works investigate LLMs in code execution (i.e. output prediction). Early studies Austin et al. (2021); Nye et al. (2021) indicated that LLMs face challenges in accurately executing code but can benefit from intermediate reasoning steps. Subsequent research has utilized various strategies to enhance execution performance, including execution-informed pretraining Liu et al. (2023a), CoT prompting with trace supervision Ni et al. (2024), and iterative instruction prompting Lyu et al. (2024). Crux Gu et al. (2024) formalized input and output prediction tasks specifically for evaluating LLMs’ code understanding and reasoning, while LCB extended these assessments to realistic programs written by humans. REVAL Chen et al. (2024b) and CACP Hooda et al. (2024) uncovered inconsistencies and concept-level misunderstandings of LLMs in code through intermediate reasoning analysis in code execution. CodeScore Dong et al. (2024), Lever Ni et al. (2023), and xCodeEval Khan et al. (2024) have leveraged code execution to improve model evaluation methodologies and enhance the performance of downstream tasks.
Robustness evaluation is another research direction in assessing LLMs for code. Prior studies focused on prompt-level variations in code generation, examining how semantically equivalent NL descriptions impact model performance Mastropaolo et al. (2023); Chen et al. (2024a); Zhuo et al. (2023). Apart from NL perturbations, researchers have investigated model robustness under programming language perturbations. For example, CCTest Li et al. (2023) introduced structural mutations, CREAM Gao et al. (2023) employed counterfactual reasoning, and ReCode Wang et al. (2022) applied semantic-preserving transformations, advancing the robustness in code comprehension and completion. BigCodeBench Zhuo et al. (2025) further explored generalization under complex function composition and tool usage. Furthermore, Chen et al. (2025); Cuadron et al. (2025) and Kumar et al. (2025) have recently revealed that LRMs frequently suffer from overthinking and produce unnecessary analysis and unstable thoughts in math-related works.
3 Methodology
This section outlines our benchmark methodology, including dataset selection, benchmark pipeline, perturbation strategies, and experimental settings.
3.1 Benchmark Tasks and Datasets
We evaluate two benchmarks: Crux Gu et al. (2024) and LCB Jain et al. (2025). Crux contains 800 synthetic problems generated by Code-LLaMA-34B, each featuring concise Python snippets with generic function names f and a corresponding input-output pair for input and output prediction tasks. In contrast, LCB contains 479 real-world coding problems sourced from 85 human-submitted LeetCode solutions, focusing solely on output prediction. While both datasets avoid excessive memory or runtime demands, LCB problems exhibit more complex structures, such as nested functions, decorators, and external libraries, requiring more computation.
@@ 3.2 Benchmark Execution Pipeline
Our benchmark follows a unified evaluation pipeline (Figure 3) across all perturbations. Each original code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to remove comments and standardize formatting, producing a clean baseline “vanilla” version (VAN). Subsequently, we apply a selected perturbation and regenerate the code using AST unparsing to maintain formatting consistency (except for comment-based perturbations). Each perturbed code snippet is executed with the provided inputs to ensure syntactic validity and functional correctness. The validated perturbed code is then provided to LLMs, and we record all execution results and evaluate them using the original evaluation methods from the dataset to compute the Pass@1 score.
@@ 3.3 Perturbation Strategies
We design two categories of perturbations to evaluate model robustness comprehensively: (1) Code Structural Perturbations, which modify code structure without affecting functionality, and (2) Textual Distraction Perturbations, which inject misleading NL cues and hints. Detailed examples and principles are provided in Appendix A.
3.3.1 Code Structural Perturbations
(1) Renaming Entities (REN).
[…]
(2) Reformatting Conditional Expressions (RTF).
We rewrite conditional expressions within “if” and “while” statements into logically equivalent but structurally complex and indirect forms. We use 16 predefined transformation templates to reformat comparison, boolean, and other conditional expressions. This perturbation evaluates the LLMs’ ability to maintain logical consistency and reasoning under increased complexity.
(3) Inserting Garbage Code (GBC).
We insert syntactically valid but functionally irrelevant “garbage” code segments to challenge model robustness against code structural distractions. Three specific types include: (1) repeated global variable declarations mirroring local identifiers, (2) non-executable conditional dead branches, and (3) uninvoked infinite recursive function definitions.
(4) Aggregating Structural Perturbations (ALL).
We comprehensively aggregate all structural perturbations (REN, RTF, GBC) together and create a highly adversarial scenario, thoroughly challenging LLM robustness across multiple structural perturbations simultaneously.
3.3.2 Textual Distraction Perturbations
(1) Inserting Misleading Descriptive Comments (MDC).
We insert misleading NL comments for eight key AST nodes (e.g., function definitions, returns, loops, conditionals) to evaluate LLM robustness against textual cues and contradictory information. We initially instruct GPT-4o to generate those messages and manually filter them to ensure they are generic, explicitly incorrect, and contradict the actual code logic. Additionally, we introduce a probability parameter $p$ to vary comment density and analyze its impact and validity on model robustness in §5.1.
(2) Inserting Misleading Print Statements (MPS).
Similar to MDC perturbation, we inject misleading NL messages as executable print statements. Unlike comments that AST-based heuristics can easily filter out, print statements are logically irrelevant to functionality and should not be perceived as semantic guides. Allowing us to directly compare the behavior with MDC and distinguish models’ NL reliance in different embedding formats. Since MDC and MPS perturbations inject shallow and contextually misleading information, we categorize them as contextual-level textual perturbations.
(3) Inserting Misleading Hint Comments (MHC).
We instruct GPT-4o to generate logically plausible but incorrect hints that contradict the actual program behavior and the ground truth. After verifying their incorrectness, they are strategically inserted into function definitions or return statements using AST parsing. Since these plausible hints challenge the model’s high-level inference ability, we classify MHC as a reasoning-level textual perturbation.
4 Experiments
This section presents a comprehensive analysis of model robustness under the perturbation strategies introduced in §3.3. We assess the reliability of LLMs across different perturbation types, model scales, reasoning modes, and prediction tasks.
@@ 4.1 Experiment Settings
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o) Hurst et al. (2024), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet) Anthropic (2024b, a), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro) Pichai and Hassabis (2024); Pichai et al. (2024), DeepSeek-V3 DeepSeek-AI et al. (2025b), Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder-Ins) Yang et al. (2025); Hui et al. (2024), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B) Grattafiori et al. (2024). Additionally, we assess three cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al. (2025a), and QwQ-32B Qwen Team (2025).

### v3

# CodeCrash: Exposing LLM Fragility to Misleading Natural Language in Code Reasoning

@@ Abstract
Large Language Models (LLMs) have recently demonstrated strong capabilities in code-related tasks, but their robustness in code reasoning under perturbations remains underexplored. We introduce CodeCrash, a stress-testing framework with 1,279 questions from CruxEval and LiveCodeBench, designed to evaluate reasoning reliability under structural perturbations and misleading natural language (NL) contexts. Through a systematic evaluation of 17 LLMs, we find that models often shortcut reasoning by over-relying on NL cues, leading to an average performance degradation of $\mathbf{23.2\%}$ in output prediction tasks. Even with Chain-of-Thought reasoning, models on average still have a $\mathbf{13.8\%}$ drop due to distractibility and rationalization, revealing a lack of critical reasoning capability to distinguish the actual code behaviors. While Large Reasoning Models with internal reasoning mechanisms improve robustness by fostering critical thinking, plausible yet incorrect hints can trigger pathological self-reflection, causing 232-3 times token consumption and even catastrophic cognitive dissonance in extreme cases for QwQ-32B. We refer to this phenomenon as Reasoning Collapse. CodeCrash provides a rigorous benchmark for evaluating robustness in code reasoning, guiding future research and development toward more reliable and resilient models.
Website and Leaderboard: https://cuhk-arise.github.io/CodeCrash/
@@ 1 Introduction
Large Language Models (LLMs), exemplified by GPT families OpenAI et al. (2024); Hurst et al. (2024) and the DeepSeek series DeepSeek-AI et al. (2025b, a), have recently exhibited remarkable performance across diverse code-related tasks, including code generation Zan et al. (2023); Liu et al. (2023b); Xiao et al. (2024), code completion Ding et al. (2023), program repair Fan et al. (2023); Liu et al. (2025), program testing Deng et al. (2023); Kang et al. (2023), and test case generation Steenhoek et al. (2025); Takerngsaksiri et al. (2025). These advancements have facilitated the incorporation of LLMs into Integrated Development Environments, such as GitHub Copilot GitHub (2022), OpenAI Codex Chen et al. (2021), and Tabnine Tabnine (2023), highlighting their potential to automate various aspects of software development Zhang et al. (2023); Pandey et al. (2024). The effectiveness of these applications heavily depends on the model’s ability to comprehend the program logic, underlying functionality, and developer intent Pan et al. (2024), but real-world codebases are often disorganized, with ambiguous identifiers Gao et al. (2023), garbage code Li et al. (2023), and inconsistent comments Rong et al. (2025). Those noises can mislead models into misinterpretations and incorrect inferences, compromising the model’s reliability in downstream tasks.
Traditional robustness studies and obfuscators have primarily focused on code structure mutations, such as renaming identifiers, modifying control flow, or inserting unreachable code Wang et al. (2022); Wan et al. (2022); Li et al. (2023). While they are effective for assessing model pattern-matching, these variations remain at the structural level. On the other hand, natural language (NL) perturbations have mainly been explored in NL-to-Code tasks (e.g., code generation and completion) by editing the task descriptions Wang et al. (2022); Zhuo et al. (2023); Mastropaolo et al. (2023); Chen et al. (2024a), which evaluates prompt sensitivity rather than robustness in code reasoning. To the best of our knowledge, CodeCrash is the first benchmark to systematically utilize NL-embedded misleading perturbations to evaluate whether LLMs can prioritize executable semantics over NL cues in code reasoning and understanding.
To diagnose whether LLMs truly understand program logic, we adopt input and output predictions from CruxEval (CruxGu et al. (2024) and extend them with LiveCodeBench (LCBJain et al. (2025) to cover real-world programs. We categorize our NL-embedded perturbations into contextual-level (obviously incorrect cues in multiple formats) and reasoning-level (plausible but incorrect hints inducing rationalization behaviors Chen et al. (2025b)), with the goal of stress-testing the critical reasoning capability of LLMs. Furthermore, building upon program structure-consistent (PSC) mutations from CCTest Li et al. (2023) and other perturbations (e.g., dead-loop poisoning) Wan et al. (2022); Sun et al. (2023), we construct an aggregated structural perturbation that integrates all traditional structural perturbations for representative comparison.
We perturb questions in Crux and LCB, and evaluate seventeen LLMs in the direct inference setting, observing significant performance degradation across all designed perturbations. We then extend our evaluation to Chain-of-Thought (CoT) Kojima et al. (2022) reasoning, and observe the phenomena of distractibility Shi et al. (2023) and rationalization Barez et al. (2025) under contextual-level and reasoning-level perturbations, respectively, exposing the insufficiency of critical thinking in semantic tracing and code understanding. Therefore, we further assess three advanced Large Reasoning Models (LRMs), which achieve superior robustness because of their enhanced internal reasoning. However, we find that LRMs are trained to be overly cautious of uncertainty: contextual-level perturbations increase reasoning tokens to digest the obviously misleading cues Kumar et al. (2025), while reasoning-level perturbations exploit this bias to trigger pathological overthinking and even “collapse” in QwQ-32B, constituting a novel failure mode. In addition, we accidentally identify a potential limitation in input prediction for evaluating code comprehension, where models exploit those misleading cues and bypass the core execution logic without understanding the code.
@@
  • We introduce a perturbation framework on code reasoning tasks that covers both structural and NL-embedded code transformation for stress-testing LLM code reasoning robustness.
  • We reveal that LLMs lack sufficient critical reasoning and exploit superficial cues to shortcut reasoning and rationalize comments, extending prior findings on rationalization to code reasoning.
  • We identify a pathological self-reflection phenomenon in LRMs and novel Reasoning Collapse failure in QwQ-32B, arising from residual rationalization that conflicts with reason-faithful objectives.
2 Functionally Equivalent Perturbations
Figure 1: Overview of the CodeCrash pipeline. Perturbed parts are visualized with color highlights.
@@ 2.1 Benchmark Execution Pipeline
CodeCrash provides a unified perturbation-evaluation pipeline, as illustrated in Figure 1. Each code snippet is initially processed via Abstract Syntax Tree (AST) parsing and unparsing to standardize formatting, producing a clean “vanilla” baseline (VAN). Subsequently, we apply a selected perturbation and regenerate the code through AST parsing, ensuring that replacement, reformatting, and insertion are precisely controllable. Each perturbed code snippet is executed with the provided inputs to double-check the syntactic validity and functional correctness. The validated and perturbed code is then provided to LLMs, where we record their generated responses, extract the answers, and execute them to compute the Pass@1 accuracy Chen et al. (2021).
@@ 2.2 Perturbation Strategies
We design a comprehensive perturbation framework that includes structural, contextual-level, and reasoning-level perturbations to systematically stress-test LLMs’ code reasoning reliability. Detailed illustrations of each perturbation are provided in Appendix B.
Aggregated Structural Perturbation (PSC-ALL).
We adapt PSC mutations from CCTest Li et al. (2023) and dead-loop poisoning Wan et al. (2022); Sun et al. (2023) by aggregating them on the identifier-level (i.e., Renaming Entities, REN), instruction-level (i.e., Reformatting Conditional Expressions, RTF), and block-level (i.e., Garbage Code, GBC; including unexecutable dead loops, dead blocks, and global variables). PSC-ALL constructs a more complicated but functionally equivalent program and represents the traditional structural perturbation that modifies syntax to expose the pattern-matching biases of LLMs. For completeness, we specify REN, RTF, and GBC in Appendix B.1.1-B.1.3, and highlight their compounding effects in Appendix B.1. Since their impact is consistently weaker than PSC-ALL, we focus our experiments on PSC-ALL.
Contextual-Level Misleading Perturbations (MCC & MPS).
These perturbations aim to mislead models by injecting shallow NL cues into the surface context without altering the code logic. We target eight critical and ordinary AST nodes (e.g., function definitions, returns, and loops) and leverage GPT-4o to generate the misleading messages, which we manually filter to ensure they are generic, explicitly incorrect, and contradictory to the code logic. Concretely, we design two variants: (1) Misleading Code Comments (MCC), where we inject messages as code comments; and (2) Misleading Print Statements (MPS), where we embed messages via executable print statements. They evaluate whether LLMs naively include the irrelevant NL cues from the program for reasoning, using different injection formats to clarify that the phenomenon is not bound to a specific format.
Reasoning-Level Misleading Perturbation (MHC).
We further design the Misleading Hint Comments (MHC) perturbation, which differs from contextual-level perturbations by providing high-level incorrect hints about program outputs. We instruct GPT-4o to generate a plausible but incorrect answer that preserves the expected type and structure but contradicts the ground truth. After verifying the incorrectness, we utilize AST parsing to strategically inject it as a comment at function definitions or return statements to make it more confusing. Unlike shallow contradictions from contextual perturbations, MHC targets the reasoning process, assessing whether the model can critically evaluate conflicting information rather than rationalize the hint and shortcut its reasoning.
3 Experimental Results and Analysis
@@ 3.1 Experiment Settings
Table 1: Relative Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference. Darker red highlights represent more severe degradation. See [Absolute Version] for the corresponding absolute degradation.
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.3$ $53.2$ $27.9\%$ $31.9\%$ $32.6\%$ $40.4\%$ $23.2\%$ $28.0\%$ $11.6\%$ $44.3\%$ $28.4\%$
GPT-4o $71.3$ $64.5$ $15.0\%$ $28.4\%$ $14.0\%$ $29.1\%$ $17.2\%$ $24.3\%$ $6.4\%$ $26.6\%$ $18.4\%$
Claude Series Claude-3.5-Haiku-20241022 $57.4$ $58.2$ $24.0\%$ $36.9\%$ $13.8\%$ $0.0\% 11$$1.2\%$ $8.0\%$ $27.5\%$ $53.3\%$ $22.1\%$
Claude-3.5-Sonnet-20241022 $71.5$ $73.8$ $14.8\%$ $34.1\%$ $8.1\%$ $9.6\%$ $8.6\%$ $10.7\%$ $14.4\%$ $43.4\%$ $16.3\%$
Gemini Series Gemini-1.5-Flash $56.2$ $44.3$ $18.4\%$ $12.7\%$ $14.8\%$ $32.2\%$ $28.5\%$ $40.4\%$ $7.3\%$ $22.2\%$ $20.8\%$
Gemini-2.0-Flash $65.0$ $66.0$ $33.2\%$ $42.1\%$ $26.8\%$ $51.9\%$ $33.1\%$ $50.3\%$ $8.3\%$ $19.5\%$ $31.2\%$
Gemini-1.5-Pro-002 $67.4$ $56.2$ $19.9\%$ $23.2\%$ $23.0\%$ $33.0\%$ $21.7\%$ $32.8\%$ $10.0\%$ $32.5\%$ $23.0\%$
DeepSeek DeepSeek-V3 $67.9$ $67.8$ $12.9\%$ $35.6\%$ $16.6\%$ $41.4\%$ $10.1\%$ $34.2\%$ $10.7\%$ $29.7\%$ $21.1\%$
LLaMA Series LLaMA-3.1-8B-Instruct $36.0$ $34.7$ $23.8\%$ $19.7\%$ $22.5\%$ $28.5\%$ $17.8\%$ $23.1\%$ $16.7\%$ $39.2\%$ $23.0\%$
LLaMA-3.1-70B-Instruct $56.1$ $44.9$ $19.6\%$ $17.2\%$ $19.2\%$ $27.9\%$ $26.7\%$ $38.8\%$ $8.4\%$ $32.2\%$ $22.4\%$
LLaMA-3.1-405B-Instruct $63.5$ $50.7$ $19.7\%$ $16.9\%$ $6.6\%$ $4.7\% 11$$1.3\%$ $19.9\%$ $6.9\%$ $25.7\%$ $14.2\%$
LLaMA-3.3-70B-Instruct $59.9$ $48.5$ $17.3\%$ $20.1\%$ $12.4\%$ $19.8\%$ $13.5\%$ $25.4\%$ $6.8\%$ $20.9\%$ $15.9\%$
Qwen Series Qwen2.5-7B-Instruct $43.3$ $41.4$ $37.9\%$ $30.9\%$ $58.0\%$ $38.3\%$ $45.2\%$ $19.8\%$ $26.9\%$ $55.6\%$ $39.8\%$
Qwen2.5-14B-Instruct $47.8$ $49.5$ $39.8\%$ $30.0\%$ $34.9\%$ $41.4\%$ $2.5\% 33$$\% 9.3-9.3\%$ $22.2\%$ $30.2\%$
Qwen2.5-32B-Instruct $60.0$ $59.6$ $19.8\%$ $34.9\%$ $18.0\%$ $\% 23.7-23.7\%$ $23.6\%$ $13.1\%$ $30.8\%$ $23.1\%$
Qwen2.5-72B-Instruct $60.1$ $54.9$ $23.3\%$ $25.2\%$ $17.0\%$ $12.7\%$ $24.1\%$ $26.2\%$ $16.0\%$ $40.6\%$ $22.4\%$
Qwen2.5-Coder-32B-Instruct $67.0$ $56.6$ $22.3\%$ $27.9\%$ $23.2\%$ $36.6\%$ $16.6\%$ $30.3\%$ $10.5\%$ $1.7\% 22$$2.3\%$
Average $22.9\%$ $27.5\%$ $21.3\%$ $29.4\%$ $21.5\%$ $27.6\%$ $12.4\%$ $33.0\%$
$\mathbf{-24.6}\%$ $\mathbf{-24.3}\%$ $\mathbf{-23.8}\%$ $\mathbf{-20.1}\%$ $\mathbf{-23.2}\%$
@@
We evaluate 17 models of varying sizes and versions, including both open-source and closed-source LLMs: GPT family (GPT-4o-Mini, GPT-4o) Hurst et al. (2024), Anthropic Claude (Claude-3.5-Haiku, 3.5-Sonnet) Anthropic (2024b, a), Google Gemini (Gemini-1.5-Flash, 2.0-Flash, 1.5-Pro) Pichai and Hassabis (2024); Pichai et al. (2024), DeepSeek-V3 DeepSeek-AI et al. (2025b), Alibaba Qwen (Qwen2.5-7B-Ins, 14B-Ins, 32B-Ins, 72B-Ins, 32B-Coder-Ins) Yang et al. (2025); Hui et al. (2024), and Meta LLaMA (LLaMA-3.1-8B-Ins, 70B-Ins, 405B-Ins, LLaMA-3.3-70B) Grattafiori et al. (2024). Additionally, we assess 3 cutting-edge LRMs: o3-mini OpenAI (2025), DeepSeek-R1 DeepSeek-AI et al. (2025a), and QwQ-32B Qwen Team (2025).